Table 4-13
The demand schedule below pertains to sandwiches demanded per week.

Price

Harry’s
Quantity
Demanded

Darby’s
Quantity
Demanded

Jake’s
Quantity
Demanded

$3

3

4

3

$5

1

2

x

Refer to Table 4-13. Suppose Harry, Darby, and Jake are the only demanders of sandwiches. Also suppose x = 2. Then

a.

the slope of Jake’s demand curve is -1/2, and the slope of the market demand curve is -5/2.

b.

the slope of Jake’s demand curve is -1/2, and the slope of the market demand curve is -2/5.

c.

the slope of Jake’s demand curve is -2, and the slope of the market demand curve is -5/2.

d.

the slope of Jake’s demand curve is -2, and the slope of the market demand curve is -2/5.
Answer #1

Slope of Harry demand curve=-2/2=-1

Darbys demand curve slope=-2/2=-1

slope of Jakes demand curve=-2/1=-2

slope of market demand curve=-2/5

ans is D